40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 19 16:20:52 2015, ericq, Update, LSC, Green beatnote couplers installed 
    Reply  Mon Oct 26 18:59:25 2015, gautam, Update, LSC, Green beatnote couplers installed 
Message ID: 11714     Entry time: Mon Oct 26 18:59:25 2015     In reply to: 11700
Author: gautam 
Type: Update 
Category: LSC 
Subject: Green beatnote couplers installed 

I found (an old) 10 dB coupler in the RF component shelves near MC2 - it has BNC connectors and not SMA connectors, but I thought it would be worth it to switch out the 20dB coupler currently on the X green beat PD on the PSL table with it. I used some BNC to SMA adaptors for this purpose. It appears that the coupler works, because I am now able to register an input signal on the X arm channel of the digital frequency counter (i.e. the coupled output from the green beat PD). I thought it may be useful to have this in place and do an IR transmissions arm scan using ALS for the X arm as well, in order to compare the results with those discussed here. However, the beat note amplitude on the analyzer in the control room looks noticeably lower - I am not sure if the coupler is responsible for this or if it has to do with the problems we have been having with the X end laser (the green transmission doesn't look glitchy on striptool though, and the transmission itself is ~0.45). In any case, we could always remove the coupler if this is hindering locking efforts tonight. 

ELOG V3.1.3-